OpenStudy (anonymous):

Express the following equation in slope-intercept form: y - 11 = - 5 (x + 8) - 2

OpenStudy (radar):

Is the expression on the right of the equal sign divided by 2?

OpenStudy (anonymous):

I'm not sure what slope-intercept form means, but... y-11 = -5(x+8) is of the form y-y1 = m(x-x1) where m is the slope. so, m, the slope would be m= y-y1/x-x1 or, -5=y-11/x+8

OpenStudy (radar):

Assuming the problem is: y-11 -5(x+8) ----- = --------- 1 8 cross multiplying getting: -5(x+8)=8y -88 -5x-40=8y-88 Isolating y: 8y=-5x -40 + 88 8y=-5x+48 y=(-5/8)x+6 In slope intercept form Note m= -5/8 (slope)